A solid iron rectangular block of dimensions 127x52x15 mm and a density of 7860 kg/ m3 is suspended in
brine solution. When the iron block is completely immersed in the solution.Determine
i) the upthrust on the Iron block
ii) tension in the supporting cable
1
Expert's answer
2018-01-09T05:03:52-0500
The volume of the block is 127 X 52 X 15 mm^3
=0.127 × 0.052 × 0.015 m^3
=9.906 × 10^-5 m^3

Again density of the block was = 1786 kg/m^3
Therefore the mass of the block will be = 9.906 ×10^-5 ×1786 kg
=0.1769 kg

Again , the density of saturated brine solution at room temperature = 1202 kg / m^3

i) The upthrust will be = volume of displaced liquid × density of the liquid × g
= 9.906 × 10^-5 × 1202 x 9.8 N
=1.167 N
ii) tension of the supporting thread is
= Weight of the block - upthrust
= 0.1769 × 9.8 - 1.167 N
= 1.734 - 1.167 N
= 0.033 N